A ball is drawn randomly from a jar that contains 3 red balls, 6 white balls, and 9 yellow balls. Find the probability of the given event, and show your answers rounded to 4 decimal places when possible.

a. A red ball is drawn. P(red) = _____

b. A white ball is drawn. P(white) = _____

c. A yellow ball or red ball is drawn. P(yellow or red) = ____
Answer #1

Solution

Total balls = 6 + 6 + 9 = 18

(a)

P(red) = 3 / 18 = 0.1667

(b)

P(white) = 6 / 18 = 0.3333

(c)

P(yellow or red) = 9 / 18 + 3 / 18 = 12 / 18 = 0.6667